Operations Report – March 22nd

Crew 207 Operations Report 22 Mar 2019

SOL: 6

Name of person filing report: Tom Baldwin

Non-nominal systems: Generator:Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator

Generator (hours run): 13 hours

Generator was turned off at 11:00, SOC 100%, and turned on at 20:00

Diesel Reading – 50 %

Propane Reading – 63 %

Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use

Water (static tank) – 180 gallons

Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 154 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 142014.9 gallons

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:53.6 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Curiosity rover used: Not used

Hours: 84.4 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A

ATV Fuel Used: N7A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where?

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts,

(including the three used for EVAs today).

Summary of Hab operations: Checked SOC of electrical system.

07:45 100%, generator on

11:55 100%, generator off

13:50 94% generator off

17:55 85% generator off

Summary of Green Hab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None